OpenStudy (anonymous):

One x-intercept for a parabola is at the point (1,0). Use the factor method to find the other x-intercept for the parabola defined by this equation: y=5x^2+5x-10 Separate the values with a comma.

OpenStudy (anonymous):

hmm, sorry im not sure about this one...

OpenStudy (anonymous):

what does ^ mean?

OpenStudy (anonymous):

^ means exponent it is a small number.

OpenStudy (anonymous):

y = −5x² + 15x − 10 y = −5 (x² − 3x + 2) y = −5 (x − 1) (x − 2) x−intercepts: (1,0), (2,0)

OpenStudy (anonymous):

It looks like it would be right but it was wrong. :( Thanks for trying.